OpenStudy (anonymous):

a passenger car will go 434 miles on 15.5 gallons of gasoline in city driving. what is the rate in miles per gallon?simplify answer.

OpenStudy (ajprincess):

Divide 434 by 15.5 and that will giv you the answer.

OpenStudy (anonymous):

but it would be a deciimal wouldnt it?

OpenStudy (anonymous):

Solve for x:\[\frac{434}{15.5}=\frac{x}{1}\]

OpenStudy (ajprincess):

nope. it will be a whole number.

OpenStudy (anonymous):

ooh i was doing it wrong, but i got a decimal still. but i was multiplying

OpenStudy (anonymous):

\[\frac{434}{15.5}=\frac{4340}{155}=28 \text{ mpg}\]

OpenStudy (anonymous):

i got 28 mpg too thanks